solve this quadratic equation
x^4-5x^2+4=0

Answer by stanbon(75887)

x^4-5x^2+4=0
---
Factor:
(x^2-4)(x^2-1) = 0
---
(x-2)(x+2)(x-1)(x+1) = 0
x = -2 or -1 or 1 or +2
===========================
If you use a "u-substitution" you get:
Let u = x^2
Then u^2-5u+4 = 0
Factor:
(u-4)(u-1) = 0
u = 4 or u = 1
---
Convert to "x" solutions:
x^2 = 4 or x^2 = 1
x = +/- 2 or x = +/-1
==========================



Cheers,
Stan H.
==========